OpenStudy (anonymous):

An airplane flew 3 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. 180 mph 200 mph 220 mph 240 mph

OpenStudy (anonymous):

@shaik0124

OpenStudy (ghazi):

(U+W) *3=S (U-W)*2=S

OpenStudy (shaik0124):

240mph